OpenStudy (anonymous):

If the sides of a square are lengthened by 5 cm, the area becomes 144c^2. Find the lengrh of a side of the original square.

OpenStudy (anonymous):

the new side length "s" is the original unknown length "x" plus 5 cm s=x+5 and the area of a square is s^2 so: (x+5)^2=144 take the square root of both sides x+5=12 x=7 so the original side length was 7cm

OpenStudy (anonymous):

thank you for your help